OpenStudy (anonymous):

Determine whether the function is linear. Then find the evaluate for the given value of x. f(x)=lx+4l;f(3) f(x)=3x-4;f(-1) f(x)=-4x;f(8)

OpenStudy (anonymous):

1.) No. f(3) = |(3)+4| = 7 2.) Yes. f(-1) = 3(-1) - 4 = -7 3.) Yes. f(8) = -4(8) = -32 You should start to see what the questions are asking.. linear means a straight line. f(number) simply means plug (number) into the equation.

OpenStudy (anonymous):

I can't tell if it's linear though, so we isn't the first one linear? Because it has absolute value?

OpenStudy (anonymous):

Right. absolute value means it will make a V shape.

OpenStudy (anonymous):

Ok, thank you so much! I finally get it! :)

OpenStudy (anonymous):

You're welcome! If you have any more questions just feel free to ask me.